Given that f(x)= x^2-x-3, find the following:
f(x+h) - f(x)
_____________
h

Answer by stanbon(75887)

Given that f(x)= x^2-x-3, find the following:
[f(x+h) - f(x)] / h

===========================
f(x+h) = (x+h)^2 - (x+h) -3
f(x+h) = x^2 + 2hx + h^2 -x-h -3
------
f(x) = x^2 - x -3
=================================
f(x+h)-f(x) = 2hx + h^2 -h
=================================
[f(x+h)-f(x)]/h = 2x + h -1
==================================
Cheers,
Stan H.
==================================